Jorge buys a property on credit, having to pay it in 4 years, in the 1st year the monthly installments are US$ 660. In successive years the monthly installments suffer an increase of 3% in relation to with those of the previous year. What is the approx. of each monthly installment in the last year of payment? a) $700
b) 7215
c) $743
d) $765

Answer by Boreal(15235)

This is 660(1.03)^3=$721.20
if b is $721.5, that would be the best answer
year 1 660
year 2 660+(660*.03)=679.80
year 3 679.80+(679.8*0.03)=700.19
year 4 700.19+(700.19*0.03)=$721.20
This is a bit tricky in that the amount rises for only 3 years, because for the first full year the monthly installments were $660. Year 2 represents the first year of a higher rate; year 4 represents the third and last year of the higher rate.